OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 5x – 9y = –16 2x + 6y = –16

jhonyy9 (jhonyy9):

multiply first with 2 and the second with 5 and you get 10x-18y=-32 10x+30y=-80 so after this subtract from first the second -------------- 0 -48y=48 y= 48/-48 = -1 y= -1 2x+6y= -16 2x+6(-1)= -16 2x -6 = -16 2x= -10 x= -10/2 = -5 x= -5 y= -1

jhonyy9 (jhonyy9):

2(-5)+6(-1)=-16 -10-6=-16 -16=-16

jhonyy9 (jhonyy9):

hope that this is understandably

jhonyy9 (jhonyy9):

if do you like it give me one medal

jhonyy9 (jhonyy9):

thnx

jhonyy9 (jhonyy9):

bye